A rock is lying at the top of a 2 m high cliff and it possesses a potential energy of 80 J. What is the mass of the rock?




1
Expert's answer
2021-02-02T19:46:24-0500

By the definition of the potential energy, we have:


"PE=mgh,""m=\dfrac{PE}{gh}=\dfrac{80\ J}{9.8\ \dfrac{m}{s^2}\cdot2\ m}=4.08\ kg."